Could someone show me how to solve the below equation step by step - I am doing something wrong and cannot get it to come out correctly where it the solution can verify:
18 - (x+3) - 2x = 4 + (x-5)
Thank you in advance

Found 4 solutions by Alan3354, greenestamps, ikleyn, Edwin McCravy:
Answer by Alan3354(69443)

Could someone show me how to solve the below equation step by step - I am doing something wrong and cannot get it to come out correctly where it the solution can verify:
18 - (x+3) - 2x = 4 + (x-5)
====================
How do you know it's wrong?
------
Show what you've tried.
Knowing what you're doing wrong could be important.

Answer by greenestamps(13200)



In your earlier post, you went from

18 - (x+3) = ...

to

18 - x + 3 = ...

That is not right.

18 - (x+3) --> 18-x-3 (not 18-x+3)

It's a basic mistake beginning algebra students make....

"18 - (x+3)" means you are subtracting BOTH "x" AND "3"; so the result is 18-x-3.


Answer by ikleyn(52778)

.

If you want to learn on how to solve simple and simplest linear equations,  read the lesson
    - HOW TO solve a linear equation
in this site.

Consider this lesson as your textbook, handbook, tutorials and (free of charge) home teacher.

Happy learning (!)



Answer by Edwin McCravy(20054)

The others did not see your earliest post.  This is what you posted
that you had done:

18 - (x+3) - 2x = 4 + (x-5)
18 - x +3 - 2x = 4 + x-5 <---THIS STEP IS WRONG!!!!!)
-x - 2x +18 + 3 = 4 + x-5
-3x +21 = 4 + x-5
-3x - x + 21 - 21 = x - x + 4 -5 -21
4x = -22
x = -5.5

------------------------------------
Your error was in removing parentheses when there is a minus sign preceding
the parentheses.

You changed  - (x+3) to -x+3 when you should have changed it to -x-3.

BOTH SIGNS are changed when removing the parentheses, not just the first
one.

When an expression in parentheses is preceded by a minus or negative sign,
it is the same as if you were distributing -1 into the parentheses.

So when you see -(x+3), consider it to be 

                -1(x+3)

then distribute like this

               (-1)(x)+(-1)(3)

then

                -x + (-3)

and finally

                -x - 3

Now do the problem again, using -x - 3, not what you had.      

Edwin